OpenStudy (anonymous):

Find the determinant of the following matrix. [1/3, 3 5/6, 3/4]

OpenStudy (earthcitizen):

-9/4

OpenStudy (earthcitizen):

\[\left[\begin{matrix}1/3 & 3 \\ 5/6 & 3/4\end{matrix}\right]\] = (1/4)-(10/4) = -9/4

OpenStudy (anonymous):

-2 1/4?

OpenStudy (earthcitizen):

determinant = ad-cb \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]

OpenStudy (anonymous):

\[\frac{1}{3}\times \frac{3}{4}-3\times \frac{5}{6}\]

OpenStudy (anonymous):

what earthcitizen said